Starting from home, you walked 200.0 m N going to the market. From there, you

proceeded 300.0 m E to a grocery store before walking250.0 m SE to buy computer

parts. You decided then to go to school which is 100.0 m 30O N of E of the computer

shop. Applying the concepts of vector addition using the graphical and component

method, what is your latest location with respect to your house? Show complete

solutions of both methods used.



1
Expert's answer
2021-11-04T10:18:05-0400

"s_x=300+250\cos 45°+100\cos 30°=563.4~m,"

"s_y=200-250\sin45°+100\sin 30°=73.2~m,"

"s=\sqrt{s_x^2+s_y^2}=568.1~m,"

"\theta=\arctan\frac{s_y}{s_x}=7.4°" N of E.